# InterActHuman: Multi-Concept Human Animation with Layout-Aligned Audio Conditions

### v1

@@ Abstract
End-to-end human animation with rich multi-modal conditions, e.g., text, image and audio has achieved remarkable advancements in recent years. However, most existing methods could only animate a single subject and inject conditions in a global manner, ignoring scenarios that multiple concepts could appears in the same video with rich human-human interactions and human-object interactions. Such global assumption prevents precise and per-identity control of multiple concepts including humans and objects, therefore hinders applications. In this work, we discard the single-entity assumption and introduce a novel framework that enforces strong, region-specific binding of conditions from modalities to each identity’s spatiotemporal footprint. Given reference images of multiple concepts, our method could automatically infer layout information by leveraging a mask predictor to match appearance cues between the denoised video and each reference appearance. Furthermore, we inject local audio condition into its corresponding region to ensure layout-aligned modality matching in a iterative manner. This design enables the high-quality generation of controllable multi-concept human-centric videos. Empirical results and ablation studies validate the effectiveness of our explicit layout control for multi-modal conditions compared to implicit counterparts and other existing methods. See our Project Page for more details.
@@ 1 Introduction
By leveraging the priors of pretrained Diffusion Transformer-based (DiT) video diffusion models bar2024lumiere ; svd ; ayl ; guo2024animatediff ; zhou2022magicvideo ; walt ; wang2023modelscope ; vdm ; videogan ; cvideogan ; singer2022make ; text2video ; villegas2022phenaki ; lin2025apt , end-to-end human animation models, especially audio-driven approaches he2023gaia ; tian2024emo ; xu2024hallo ; wang2024vexpress ; chen2024echomimic ; xu2024vasa ; stypulkowski2024diffused ; jiang2024loopy ; lin2024cyberhost ; lin2025omnihuman has achieved high-quality human-centric video generation and strong controllability from multi-modal conditions, such as text, image and audio. However, most of existing methods commonly hold an assumption of a single-identity paradigm: all available conditions should be fused globally and implicitly assumed to describe one unique subject in the given image. Although this global injection strategy simplifies conditioning by sharing the same condition signal across all regions, it fundamentally limits scalability in scenarios involving multiple individuals or complex human-object interactions, where each entity requires distinct appearance and voice attributes.
@@
Recent multi-concept video customization methods, such as Video-Alchemist chen2025multi , ConceptMaster huang2025conceptmaster , Phantom liu2025phantom , and SkyReels-A2 fei2025skyreels , enable injecting multiple reference images into a single video, facilitating multi-person or human-object interaction scenarios. However, these methods face significant challenges when directly applied to human animation tasks. In such tasks, animation signals are highly specific to individual identities and demand precise conditioning and alignment with specific spatiotemporal regions. For example, audio signals are exclusively associated with the current speaker and are unrelated to the background or other individuals’ concepts. Existing customization methods, akin to end-to-end human animation approaches, still adopt video-level condition injection. While this approach may suffice for general video generation, it often causes confusion in human-centric video generation, making it difficult to produce satisfactory results, as shown in Fig. 3. For end-to-end human-centric video generation, conditioning inputs should ideally include not only global modalities (e.g., reference images and text descriptions) but also local modalities (e.g., audio). As mentioned, existing human animation and multi-concept customization methods fail to address this critical distinction. This limitation motivates us to propose a new framework that enables the precise injection of local human-related modalities, a capability that is both essential and urgently needed for robust multi-concept, human-centric video generation, as illustrated in Fig. 1.
@@
In this paper, we propose InterActHuman, a video diffusion framework to achieve spatial alignment of multi-modal conditions for multi-concept human video generation. Unlike previous methods chen2025multi ; huang2025conceptmaster , which rely on feature fusion and attention mechanisms to implicitly learn the relationships between condition signals and concepts, InterActHuman introduces an attention module that explicitly predicts the spatial locations where reference concepts appear in the video. This explicit alignment allows the model to accurately associate local audio conditions with the correct spatial regions via iterative mask prediction and masked audio attention during the diffusion inference process. Our framework offers two key advantages. First, it enforces a stronger layout constraint by explicitly and precisely binding conditions to their corresponding spatial regions. Second, it provides a unified interface for synchronously injecting all modalities such as visual and acoustic inputs, through the layout. These features make InterActHuman well-suited for multi-modal multi-concept human animation and serve as a baseline framework for this domain.
@@
Although using an explicit mask for condition injection might seem straightforward at first glance, it introduces an intrinsic chicken-and-egg dilemma: during inference, the final video is not yet available, leaving the spatial positions of each identity uncertain, thus making accurate mask prediction impossible; yet without those masks, the spatial injection of local audio conditions cannot be achieved, resulting in an incomplete or misaligned video generation process. To address this dilemma, we leverage the iterative denoising process inherent in diffusion models song2021ddim . Specifically, we introduce a mask predictor branch in the diffusion pipeline and adopt an interleaved mask prediction strategy, wherein the mask predictions generated at step $k$ serve as guidance for condition injection in step $k+1$. This iterative refinement breaks the cyclic dependency by progressively ‘discovering’ the spatial layout, ensuring that even in the absence of ground-truth video during inference, precise spatial alignment is achieved. In essence, our approach transforms the chicken-and-egg problem into a sequential and converging process that robustly aligns local audio conditions.
@@ 2 Related Works
Video Diffusion Models have enabled unprecedented quality text-to-video (T2V) or image-to-video (I2V) generation in recent years, thanks to diffusion-based generative models jonathan2020ddpm ; song2021ddim ; karras2022edm ; song2020score ; liu2022flow . Early T2V approaches either adapt pretrained text-to-image networks in a training-free manner singer2022make ; wu2023tune ; qi2023fatezero or fine-tune UNet-based latent diffusion architectures guo2024animatediff ; zhou2022magicvideo ; svd ; wang2023modelscope . To push the frontier, recent works compress spatiotemporal features with 3D causal VAEs yu20233DVAE , and migrate to Diffusion Transformer (DiT) backbones vaswani2017attention ; peebles2023scalable ; videoworldsimulators2024 ; hong2022cogvideo . Through progressive low-to-high resolution pretraining and fine-tuning polyak2024moviegen ; kong2024hunyuanvideo ; wang2025wan , these models yield longer, more coherent, and high-quality videos. Our method is also built upon the pretrained DiT video generation models.
@@
Human Animation Models synthesize videos of people driven by text, reference images, human body poses or audios. Early GAN-based methods siarohin2019fomm ; zhao2022tps ; siarohin2021mraa ; jiang2024mobileportrait ; wang2021facev2v are trained on small datasets nagrani2017voxceleb ; siarohin2019fomm ; xie2022vfhq ; zhu2022celebv for self-supervised pose transfer. Diffusion-based approaches shao2024human4dit ; zhang2024mimicmotion ; aa ; DBLP:conf/nips/00010ZF0LTCX0L24 ; wang2025multi now surpass GAN-based methods by conditioning on 2D skeletons, 3D depth, or mesh sequences. Audio-driven portrait methods GeneFace ; zhang2023sadtalker ; EMO ; jiang2024loopy ; hallo3 have been extended toward full-body motion via two-stage pipelines for improved hand quality VLogger ; EchomimicV2 ; EMO2 ; diffted and one-stage unified framework designs lin2024cyberhost ; lin2025omnihuman . In summary, none of them explored multi-concept human animation, and our method is the first to enable multi-concept human animation with local audio conditions.
@@
Multi-Concept Video Customization Models have received limited attention. Early single-concept identity-preserving methods include Videobooth jiang2024videobooth , which learns coarse-to-fine embeddings from WebVid bain2021frozen via Grounded-SAM kirillov2023segment ; liu2023grounding ; ren2024grounded ; ID-Animator he2024id , which integrates IP-Adapter ye2023ip into AnimateDiff guo2024animatediff ; and ConsisID yuan2024identity , which decouples frequency signals to preserve facial identity. More recently, Video-Alchemist chen2025multi , ConceptMaster huang2025conceptmaster , and Phantom liu2025phantom support multiple reference images and text descriptions via cross- or self-attention injection for general-purpose multi-concept customization. BlobGen-Vid feng2025blobgen further enables local layout control of text-specified concepts using user- or LLM-provided spatiotemporal masks. However, all existing methods rely solely on image and text conditions and lack support for multimodal inputs such as audio, which we argue are essential for truly versatile, human-centric video generation.
@@ 3.1 Problem Setting and Preliminaries
Preliminaries. Transformer-based text-to-video latent diffusion models have recently demonstrated remarkable capabilities in generating high-quality video content. Our proposed InterActHuman framework is built upon the MMDiT-based video generation model dit ; esser2024sd3 ; seawead2025seaweed and utilize a 3D Variational Autoencoder (VAE) kingma2013auto , which compresses input videos into a compact latent space. For training, we adopt the flow matching objective lipman2022flow , which formulates the generative process as a probability flow ordinary differential equation (ODE). This ODE transports clean latent representations $z_{0}$ to their noisy counterparts $z_{t}$ along a linear path, defined as $z_{t}=(1-t)z_{0}+t\epsilon$ at timestep $t$, where $\epsilon$ is sampled from a standard Gaussian distribution. In our setting, which incorporates multi-modal conditions (image and audio), the output of the diffusion transformer is parameterized as $v_{\Theta}(z_{t},t,c_{img},c_{audio})$. This output is supervised to predict the velocity $(z_{1}-z_{0})$, resulting in the following training objective:
$\mathcal{L}=\mathbb{E}_{t,z_{0},\epsilon}\left||v_{\Theta}(z_{t},t,c_{img},c_{audio})-(z_{1}-z_{0})\right||_{2}^{2}.$
@@
Figure 2: Illustration of our framework, which adaptively predicts masks as the spatial guidance of audio condition injection. In training, we train the mask predictor (cross-attn w/ MLP) with mask loss; in inference, we collect mask predictions to cache and leverage masks predicted from the last denoising step ($t-1$) to guide the audio cross-attn in the current denoising step ($t$).
3.2 Multi-Concept Reference Image Injection
To handle multiple appearance images, we inject appearance cues via self-attention in the original DiT layers. Each reference image $X_{i}$ is encoded into latent tensors ${\mathbf{x}_{i}}$ using the same VAE as for noisy video frames. The reference latents are then flattened into token sequences and processed by the DiT along with the noisy latents $\mathbf{v}$. The reference latents will reuse the parameters of DiT to extract features and interact with noisy latents at the self-attention layer in every DiT block, allowing appearance cues to propagate to the denoising pathway. Notably, no extra networks or additional parameters are required, thus preserving the model’s efficiency.
@@ 3.3 Layout Prediction and Local Condition Injection
Although reference image injection has been addressed, leveraging spatiotemporal layouts for local conditions like precise audio alignment remains challenging. There is a conflict between mask prediction and audio injection: if the forward pass is incomplete, masks are unavailable; if it is complete, it’s too late for local condition injection. We resolve this by using the multi-step inference process of diffusion models song2021ddim , where masks predicted at step $k$ guide multi-modal condition injection at step $k+1$. As our network learns spatiotemporal layouts for each identity, we obtain precise masks without requiring user annotations.
@@
Mask Predictor. Within each DiT block, we predict a spatio-temporal mask that quantifies how strongly each reference image should influence each video frame. We attach a lightweight mask-predictor head to each of the $L$ transformer layers. The head comprises: (1) a shared linear projection that maps the hidden video features $\mathbf{h}^{v}$ and the hidden reference features $\mathbf{h}^{r}_{i}$ to query, key, and value tensors, (2) a LayerNorm, (3) 3D RoPE positional encoding, (4) a cross-attention module, and (5) a two-layer MLP. After normalization and RoPE, we let the video tokens attend to a single reference by computing $a_{i}^{(l)}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{v}{\mathbf{K}^{r}_{i}}^{\top}}{\sqrt{d}}\right)\mathbf{V}^{r}_{i}$, where $d$ is the head dimension and $l$ indexes the current layer. The attended feature $\mathbf{a}^{(l)}_{i}$ is subsequently transformed by the MLP and passed through a sigmoid to yield a layer-specific mask $\mathbf{m}^{(l)}_{i}\in[0,1]^{T}$ for reference $X_{i}$. Then we average the mask prediction from last few DiT blocks to form the final mask. Notably, the mask predictor is always trained to predict complete region of a human, regardless of whether the reference image contains only the upper body, face, or full body of a person. This design simplifies the mask prediction task and improves consistency for the conditioning module, ensuring robust and reliable conditioning across diverse scenarios. Moreover, since the mask predictors reuse in-block features and share parameters across references, they introduce minimal computational overhead while enabling explicit spatiotemporal layout control for each identity.
@@
Mask Prediction by Caching During Inference.During inference, mask prediction requires the integration of results across all layers, which poses challenges in the early stages of denoising. At these initial steps, the latent representations of reference concepts are often too ambiguous, preventing the conditioning module from generating precise mask layouts. To address this issue, we adopt an iterative mask prediction strategy. At each denoising step, the mask predicted in the previous step is cached as a layout guide for the current step, enabling the model to progressively improve the accuracy of spatiotemporal localization. This iterative process ensures that the masks dynamically adapt to the evolving video content, maintaining consistency for each concept throughout the sequence and improving the overall conditioning quality.
@@
Local Audio Conditioning. To prepare our model for multi-concept human-centric video generation, we first pretrain it on single-identity audio-conditioned animation by adding cross-attention-based audio conditioning without masks and a mixed-conditions training strategy, following recent work lin2025omnihuman . It utilizes a new cross-attention layer to inject wav2vec baevski2020wav2vec2 audio features into each DiT block after the MMDiT layer. In multi-concept setting, we find that utilizing pretrained audio cross-attention has already lead to satisfactory per-identity audio-driven performance. So we implement local audio conditioning mainly in the iterative process in inference. Instead of processing the entire set of noisy video tokens as in previous works lin2025omnihuman ; lin2024cyberhost , we inject wav2vec features into the tokens corresponding to a specific identity, using the mask predicted in the previous step as guidance. To ensure smooth transitions in noisy video features and final output videos, we apply a linear combination of meaningful audio features and muted audio features for each noisy video token, based on the predicted mask confidence along the mask edges.
Moreover, such local audio conditioning enables us to generate multi-person videos with dialogues. As long as the dialogue audio are separated into individual audios as the input, our model could generate realistic multi-person videos where people take turns to speak, facilitating many applications.
@@
Training Loss and Strategies. The overall training loss combines a flow matching diffusion loss (see Sec. 3.1) with a focal losslin2017focal for mask classification, which mitigates the training instability and gradient explosion encountered with binary cross-entropy loss, likely due to foreground-background imbalance or low-quality masks. A frame alignment flag is employed to exclude frames with invalid or low-quality masks from loss computation, ensuring robust temporal supervision. To address the ‘copy-paste behavior’ of diffusion models, which often replicate subjects from reference images with minimal variation in camera angle or posture, we apply random masking to reference images, selectively revealing the head, full body, or clothing of the original subject appearance to encourage appearance diversity. We use an appearance conditioning ratio of face to full body of $0.7\!:\!0.3$, while balancing the contributions of the diffusion loss and focal loss equally with a ratio of $1\!:\!1$.
@@
Inference Strategies. Given $N$ reference images and a text prompt, our method use Qwen2.5-VL bai2025qwen2 as the re-phraser to get detailed descriptions from each reference images, and then integrate them with the original text prompt to form the final prompt. We utilize a shared classifier-free guidance (CFG) across audio and text. To achieve local control, audio conditions are injected solely within the predicted mask regions for each reference. Notably, early denoising steps yield unreliable masks, and using them for condition injection can suppress other references. Thus, mask-based condition injection is omitted for the first 10 denoising steps and applied subsequently using the mask predicted from the previous step. We only inject audio signals with masks into positive result while applying CFG, and use a CFG strength as 6.5 and 50 denoising steps.

### v2

@@ Abstract
End-to-end human animation with rich multi-modal conditions, e.g., text, image and audio has achieved remarkable advancements in recent years. However, most existing methods could only animate a single subject and inject conditions in a global manner, ignoring scenarios where multiple concepts could appear in the same video with rich human-human interactions and human-object interactions. Such a global assumption prevents precise and per-identity control of multiple concepts including humans and objects, therefore hinders applications. In this work, we discard the single-entity assumption and introduce a novel framework that enforces strong, region‑specific binding of conditions from modalities to each identity’s spatiotemporal footprint. Given reference images of multiple concepts, our method could automatically infer layout information by leveraging a mask predictor to match appearance cues between the denoised video and each reference appearance. Furthermore, we inject local audio condition into its corresponding region to ensure layout-aligned modality matching in an iterative manner. This design enables the high-quality generation of human dialogue videos between two to three people or video customization from multiple reference images. Empirical results and ablation studies validate the effectiveness of our explicit layout control for multi-modal conditions compared to implicit counterparts and other existing methods. Video demos are available at https://zhenzhiwang.github.io/interacthuman/
@@ 1 Introduction
By leveraging the priors of pretrained Diffusion Transformer-based (DiT) video diffusion models (Bar-Tal et al., 2024; Blattmann et al., 2023a; b; Guo et al., 2023; Zhou et al., 2022; Gupta et al., 2023; Wang et al., 2023; Ho et al., 2022; Brooks et al., 2022; Wang et al., 2020; Singer et al., 2022; Li et al., 2018; Villegas et al., 2022; Lin et al., 2025b), end-to-end human animation models, especially audio-driven approaches (He et al., 2023; Tian et al., 2025b; Xu et al., 2024a; Wang et al., 2024a; Chen et al., 2024; Xu et al., 2024b; Stypulkowski et al., 2024; Jiang et al., 2024a; Lin et al., 2024; 2025a) have achieved high-quality human-centric video generation and strong controllability from multi-modal conditions, such as text, image and audio. However, most existing methods commonly hold an assumption of a single-identity paradigm: all available conditions should be fused globally and implicitly assumed to describe one unique subject in the given image. Although this global injection strategy simplifies conditioning by sharing the same condition signal across all regions, it fundamentally limits scalability in scenarios involving multiple individuals or complex human-object interactions, where each entity requires distinct appearance and voice attributes.
@@
Recent multi-concept video customization methods, such as Video-Alchemist (Chen et al., 2025a), ConceptMaster (Huang et al., 2025), Phantom (Liu et al., 2025), and SkyReels-A2 (Fei et al., 2025a), enable injecting multiple reference images into a single video, facilitating multi-person or human-object interaction scenarios. However, these methods face significant challenges when directly applied to human animation tasks. In such tasks, animation signals are highly specific to individual identities and demand precise conditioning and alignment with specific spatiotemporal regions. For example, audio signals are exclusively associated with the current speaker and are unrelated to the background or other individuals’ concepts. Existing customization methods, akin to end-to-end human animation approaches, still adopt video-level condition injection. While this approach may suffice for general video generation, it often causes confusion in human-centric video generation, making it difficult to produce satisfactory results, as shown in Fig. 3. For end-to-end human-centric video generation, conditioning inputs should ideally include not only global modalities (e.g., reference images and text descriptions) but also local modalities (e.g., audio). As mentioned, existing human animation and multi-concept customization methods fail to address this critical distinction. This limitation motivates us to propose a new framework that enables the precise injection of local human-related modalities, a capability that is both essential and urgently needed for robust multi-concept, human-centric video generation, as illustrated in Fig. 1.
@@
In this paper, we propose InterActHuman, a video diffusion framework for spatially aligning multi-modal conditions in multi-concept human video generation. Unlike prior methods (Chen et al., 2025a; Huang et al., 2025) that rely on feature fusion and attention to implicitly learn relationships between condition signals and concepts, InterActHuman introduces an attention module that explicitly predicts the spatial locations where the reference concepts appear in the video. This explicit layout allows the model to accurately associate local audio conditions with the correct regions via iterative mask prediction and masked audio attention during diffusion inference. Our framework offers two key advantages. First, it enforces a stronger layout constraint by precisely binding each condition to its corresponding spatial region. Second, it provides a unified interface for synchronously injecting all modalities (e.g., visual and acoustic inputs) through the layout. These features make InterActHuman well suited for multi-modal, multi-concept human animation and establish a baseline for this domain.
@@
Although using an explicit mask for condition injection may seem straightforward, it creates a chicken-and-egg dilemma: during inference the final video is not yet available, leaving the spatial positions of each identity uncertain and making accurate mask prediction impossible; yet without those masks, spatial injection of local audio conditions cannot be performed, leading to an incomplete or misaligned generation process. To address this, we leverage the iterative denoising process inherent in diffusion models (Song et al., 2021). Specifically, we introduce a mask-predictor branch into the diffusion pipeline and adopt an interleaved mask-prediction strategy, wherein the mask predicted at step $k$ guides condition injection at step k+1k{+}1. This iterative refinement progressively finalizes the spatial layout, breaking the cyclic dependency and enabling precise spatial alignment even without ground-truth video during inference. In essence, our approach converts the chicken-and-egg problem into a sequential, convergent procedure that robustly aligns local audio conditions.
@@ 2 Related Works
Video Diffusion Models have enabled unprecedented quality text-to-video (T2V) or image-to-video (I2V) generation in recent years, thanks to diffusion-based generative models (Ho et al., 2020; Song et al., 2021; Karras et al., 2022; Song et al., 2020; Liu et al., 2022). Early T2V approaches either adapt pretrained text-to-image networks in a training-free manner (Singer et al., 2022; Wu et al., 2023b; Qi et al., 2023) or fine-tune UNet-based latent diffusion architectures (Guo et al., 2023; Zhou et al., 2022; Blattmann et al., 2023a; Wang et al., 2023). To push the frontier, recent works compress spatiotemporal features with 3D causal VAEs (Yu et al., 2023), and migrate to Diffusion Transformer (DiT) backbones (Vaswani et al., 2017; Peebles & Xie, 2023; Brooks et al., 2024; Hong et al., 2022). Through progressive low-to-high resolution pretraining and fine-tuning (Polyak et al., 2024; Kong et al., 2024; Wang et al., 2025a), these models yield longer, more coherent, and high-quality videos. Our method is also built upon the pretrained DiT video generation models.
@@
Human Animation Models synthesize videos of people driven by text, reference images, human body poses or audios. Early GAN-based methods (Siarohin et al., 2019; Zhao & Zhang, 2022; Siarohin et al., 2021; Jiang et al., 2024b; Wang et al., 2021) are trained on small datasets (Nagrani et al., 2017; Siarohin et al., 2019; Xie et al., 2022; Zhu et al., 2022) for self-supervised pose transfer. Diffusion-based approaches (Shao et al., 2024; Zhang et al., 2024; Hu, 2024; Wang et al., 2024c; 2025b) now surpass GAN-based methods by conditioning on 2D skeletons, 3D depth, or mesh sequences. Audio-driven portrait methods (Ye et al., 2022; Zhang et al., 2023; Tian et al., 2025b; Jiang et al., 2024a; Cui et al., 2024) have been extended toward full-body motion via two-stage pipelines for improved hand quality (Corona et al., 2024; Meng et al., 2024; Tian et al., 2025a; Hogue et al., 2024) and one-stage unified framework designs (Lin et al., 2024; 2025a). In summary, none of them have explored multi-concept human animation, and our method is the first to enable multi-concept human animation with local audio conditions.
@@
Multi-Concept Video Customization Models have received limited attention. Early single-concept identity-preserving methods include Videobooth (Jiang et al., 2024c), which learns coarse-to-fine embeddings from WebVid (Bain et al., 2021) via Grounded-SAM (Kirillov et al., 2023; Liu et al., 2023; Ren et al., 2024); ID-Animator (He et al., 2024), which integrates IP-Adapter (Ye et al., 2023) into AnimateDiff (Guo et al., 2023); and ConsisID (Yuan et al., 2024), which decouples frequency signals to preserve facial identity. More recently, Video-Alchemist (Chen et al., 2025a), ConceptMaster (Huang et al., 2025), and Phantom (Liu et al., 2025) support multiple reference images and text descriptions via cross- or self-attention injection for general-purpose multi-concept customization. BlobGen-Vid (Feng et al., 2025) further enables local layout control of text-specified concepts using user- or LLM-provided spatiotemporal masks. Ingredients (Fei et al., 2025b) proposes to predict the layout of identities in multi-image customization, yet it does not consider the audio condition and does not support multi-person talking generation. All existing methods rely solely on image and text conditions and lack support for multimodal inputs such as audio, which we argue are essential for truly versatile, human-centric video generation.
@@ 3.1 Preliminaries
Preliminaries. Transformer-based text-to-video latent diffusion models have recently demonstrated remarkable capabilities in generating high-quality video content. Our proposed InterActHuman framework is built upon the MMDiT-based video generation model (Peebles & Xie, 2023; Esser et al., 2024; Seawead et al., 2025) and utilizes a 3D Variational Autoencoder (VAE) (Kingma & Welling, 2013), which compresses input videos into a compact latent space. For training, we adopt the flow matching objective (Lipman et al., 2022), which formulates the generative process as a probability flow ordinary differential equation (ODE). This ODE transports clean latent representations $z_{0}$ to their noisy counterparts $z_{t}$ along a linear path, defined as $z_{t}=(1-t)z_{0}+t\epsilon$ at timestep $t$, where $\epsilon$ is sampled from a standard Gaussian distribution. In our setting, which incorporates multi-modal conditions (image and audio), the output of the diffusion transformer is parameterized as $v_{\Theta}(z_{t},t,c_{img},c_{audio})$. This output is supervised to predict the velocity $(z_{1}-z_{0})$, resulting in the following training objective:
[…]
@@
Figure 2: Illustration of our framework, which adaptively predicts masks as the spatial guidance of audio condition injection. In training, we train the mask predictor (cross-attn w/ MLP) with mask loss; in inference, we collect mask predictions to cache and leverage masks predicted from the last denoising step (t1t-1) to guide the audio cross-attn in the current denoising step ($t$).
Multi-Concept Reference Image Injection. To handle multiple appearance images, we inject appearance cues via self-attention in the original DiT layers. Each reference image $X_{i}$ is encoded into latent tensors $\mathbf{x}_{i}$} using the same VAE as for noisy video frames. The reference latents are then flattened into token sequences and processed by the DiT along with the noisy latents $\mathbf{v}$. The reference latents will reuse the parameters of DiT to extract features and interact with noisy latents at the self-attention layer in every DiT block, allowing appearance cues to propagate to the denoising pathway. Notably, no extra networks or additional parameters are required, thus preserving the model’s efficiency. Please refer to Appendix D for more details.
@@ 3.2 Layout Prediction and Local Condition Injection
While reference-image injection is well studied, leveraging spatiotemporal layouts for local conditions, such as precise audio alignment, remains challenging. Mask prediction and audio injection are coupled: if a denoising step is incomplete, the masks are unavailable; if it is complete, it is too late to inject local conditions. We resolve this with the multi-step inference of diffusion models (Song et al., 2021): masks predicted at step $k$ guide multi-modal condition injection at step k+1k{+}1. As the network learns spatiotemporal layouts for each identity, precise masks emerge without user annotations.
@@
Mask Predictor. In each DiT block, we predict a spatiotemporal mask quantifying how strongly each reference image should influence each video frame. A lightweight mask-predictor head is attached to each of the $L$ transformer layers, consisting of: (1) a shared linear projection that maps hidden video features $\mathbf{h}^{v}$ and hidden reference features $\mathbf{h}^{r}_{i}$ to query, key, and value tensors; (2) LayerNorm; (3) 3D RoPE positional encoding; (4) a cross-attention module; and (5) a two-layer MLP. After normalization and RoPE, video tokens attend to a single reference via $a_{i}^{(l)}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{v}{\mathbf{K}^{r}_{i}}^{\top}}{\sqrt{d}}\right)\mathbf{V}^{r}_{i}$, where $d$ is the head dimension and $l$ indexes the current layer. The attended feature $\mathbf{a}^{(l)}_{i}$ is transformed by the MLP and passed through a sigmoid to yield a layer-specific mask $\mathbf{m}^{(l)}_{i}\in[0,1]^{T}$ for reference $X_{i}$. We then average the predictions from the last few DiT blocks to obtain the final mask. Notably, the mask predictor is trained to recover the complete human region, regardless of whether the reference image shows only the upper body, face, or full body. This simplifies mask prediction and stabilizes the conditioning module, yielding robust behavior across diverse scenarios. Because the predictors reuse in-block features and share parameters across references, they add minimal overhead while enabling explicit layout control for each identity.
@@
Mask Prediction by Caching During Inference. During inference, aggregating mask signals across layers is difficult in early denoising steps, when reference concept latents are still ambiguous. We adopt an iterative strategy: at each step, the mask predicted at the previous step is cached and used as a layout prior for the current step. This progressively sharpens spatiotemporal localization, allowing masks to adapt to the evolving content and remain consistent per concept over time, thereby improving overall conditioning quality.
@@
Local Audio Conditioning. To prepare for multi-concept, human-centric video generation, we first pretrain on single-identity audio-conditioned animation by adding cross-attention–based audio conditioning (without masks) and a mixed-conditions training strategy, following (Lin et al., 2025a). Concretely, a new cross-attention layer injects wav2vec (Baevski et al., 2020) audio features into each DiT block after the MMDiT layer. In the multi-concept setting, this pretrained audio cross-attention already provides strong per-identity audio control. We therefore implement local audio conditioning primarily at inference: instead of updating all noisy video tokens (Lin et al., 2025a; 2024), we inject wav2vec features only into tokens assigned to a given identity, using the previous-step mask as guidance. To ensure smooth transitions in latent features and in the final video, we blend meaningful and muted audio features per token using the mask confidence, with soft weighting near mask boundaries. This local conditioning also enables multi-person dialogues. Given per-speaker audio tracks as input, the model generates realistic interactions in which speakers take turns, supporting a wide range of applications.
@@
Training Loss and Strategies. The overall objective combines a flow-matching diffusion loss (Sec. 3.1) with a focal loss (Lin et al., 2017) for mask classification, which stabilizes training compared to binary cross-entropy, likely due to foreground–background imbalance and occasional low-quality masks. A frame-alignment flag excludes frames with invalid or low-quality masks from loss computation, strengthening temporal supervision. To mitigate the common “copy–paste” behavior of diffusion models—replicating subjects from references with little variation in pose or viewpoint—we randomly mask reference images to reveal only the head, full body, or clothing, promoting appearance diversity. We use a face-to–full-body appearance conditioning ratio of $0.7:0.3$ and weight the diffusion and focal losses equally ($1:1$).
@@
Inference Strategies. Given $N$ reference images and a text prompt, our method uses Qwen2.5-VL (Bai et al., 2025) as a rephraser to extract detailed descriptions from each reference image and integrate them with the original prompt. We apply a shared classifier-free guidance (CFG) across audio and text. For local control, audio conditioning is applied only within the predicted mask regions for each reference. Because early denoising steps yield unreliable masks that may suppress other references, we disable mask-based injection for the first 10 steps, then enable it using the previous-step mask. We inject masked audio only on the conditional (positive) branch during CFG sampling, with a CFG scale of 6.5 and 50 denoising steps.